# Improved Robustness of Deep Reinforcement Learning for Control of Time-Varying Systems by Bounded Extremum Seeking

### v1

@@ Abstract
In this paper, we study the use of robust model-independent bounded extremum seeking (ES) feedback control to improve the robustness of deep reinforcement learning (DRL) controllers for a class of nonlinear time-varying systems. DRL has the potential to learn from large datasets to quickly control or optimize the outputs of many-parameter systems, but its performance degrades catastrophically when the system model changes rapidly over time. Bounded ES can handle time-varying systems with unknown control directions, but its convergence speed slows down as the number of tuned parameters increases and, like all local adaptive methods, it can get stuck in local minima. We demonstrate that together, DRL and bounded ES result in a hybrid controller whose performance exceeds the sum of its parts with DRL taking advantage of historical data to learn how to quickly control a many-parameter system to a desired setpoint while bounded ES ensures its robustness to time variations. We present a numerical study of a general time-varying system and a combined ES-DRL controller for automatic tuning of the Low Energy Beam Transport section at the Los Alamos Neutron Science Center linear particle accelerator.
@@ I Introduction
Deep reinforcement learning (DRL) combines aspects of optimal control theory with deep learning techniques. In DRL, deep neural networks parameterize the policy (controller), the value function, or both. DRL is based on dynamic programming, introduced by Bellman to solve sequential decision problems through the Bellman optimality principle [1]. Classical DP assumes known analytic models of the system dynamics and the reward function, which enables the synthesis of controllers $u$ that maximize cumulative return. Reinforcement learning (RL) relaxes this modeling requirement by learning from data collected as state, action, and reward transitions when the dynamics are unknown [2, 3]. DRL extends RL to high-dimensional state and action spaces by using deep neural networks to approximate value functions and/or policies. This enables the application of RL to complex systems such as image-based perception and continuous control, where tabular methods become intractable [4]. A landmark example is deep Q-learning, which approximates the optimal action value function that maximizes cumulative reward [5]. DRL is being actively investigated for a broad set of applications, including robotic control [6], particle accelerator tuning [7], and the fine-tuning of deep neural networks [8].
@@
In contrast to DRL, the field of control theory has a long history of developing model-independent feedback control algorithms for unknown time-varying systems [16, 17, 18, 19]. For systems of the form
@@
Bounded ES is a new form of ES which has a major advantage of guaranteed bounds on control efforts and parameter update rates despite acting on noisy and analytically unknown and time-varying systems [22]. It has been studied for a wide range of dynamic systems [23], and a general weak-limit averaging analysis has extended its use with non-differentiable dithering functions as well as non-periodic time-varying systems [24]. The ability to stabilize time-varying systems with guaranteed bounds on control effort has made bounded ES useful for safe hardware implementation on high energy systems, where abrupt parameter changes can easily result in damage, such as high energy charged particle accelerators [25, 26]. The bounded ES method has also been applied for GPS-denied source localization [27, 28], optimized path tracking in robotics [29], for biology-inspired 3D source seeking [30], and for tokamak stabilization [31].
@@
While DRL is a powerful method for data-based learning of controllers for dynamic systems, one of the fundamental challenges is handling unknown and rapidly time-varying systems. On the other hand, while adaptive methods such as bounded ES are robust for time-varying systems, they are inherently local feedback-based schemes that do not exploit trajectory histories, are subject to suboptimal solutions, and can converge slowly in high-dimensional parameter spaces. This paper proposes a hybrid framework that combines DRL with bounded ES to leverage the strengths of both. We train a DRL policy using large datasets to find solutions in very few steps when the test dynamics are close to the training distribution. Crucially, the ES layer is warm-started from the DRL policy seeded with DRL-recommended control parameters, which reduces transients and accelerates adaptation when conditions drift. If the system starts to quickly change with time, and the learned DRL policy is no longer valid, robust bounded ES takes over and prevents the system’s performance from severe degradation. We demonstrate our results with numerical studies of general time-varying dynamic systems as well as detailed simulation studies of a particle accelerator application with a time-varying magnetic lattice, mimicking natural accelerator behavior.
@@ II-A Bounded Extremum Seeking for Time-Varying Systems
We briefly recall the results that summarize bounded ES as developed in [22, 23, 24]. For $x\in\mathbb{R}^{n}$, consider a system of the form
@@ II-B Deep Reinforcement Learning for Feedback Control
We model the DRL as a discounted Markov decision process (MDP) $M=(S,A,p,r,\gamma)$ with $\gamma\in[0,1)$. At time $t$, the agent observes $s_{t}\!\in\!S$, applies a continuous control $a_{t}\!\in\!A$, receives $r_{t}=r(s_{t},a_{t})$, and transitions to $s_{t+1}\!\sim\!p(\cdot\,|\,s_{t},a_{t})$. The objective is
@@ II-C Motivating Example of a Time Varying System
The averaged dynamics control direction is no longer unknown because $b^{2}\geq 0$ and therefore sufficiently large $k\alpha>0$ ensures a gradient ascent of $V(\overline{x})$. As detailed in the bounded ES references, $\omega$ must be chosen sufficiently large relative to $f$. While trivial for bounded ES, this problem is extremely difficult for DRL as shown in Fig 1.
@@
III Particle Accelerator Problem formulation
Large particle accelerators are intrinsically time-varying; they are composed of thousands of coupled electromagnetic components whose characteristics drift with temperature, usage, and maintenance; diagnostics are limited and noisy, and they require continual manual retuning. The amplitude and phase of the resonant radio frequency (RF) accelerating cavities drift with time due to temperature, amplifier/distribution effects, and magnet power-supplies ripple and hysteresis break magnetic–field repeatability. Such time-varying behavior is documented for the kilometer-long Los Alamos Neutron Science Center (LANSCE) linear accelerator, for which adaptive ES methods have been developed for beam loss minimization [37]. Because ES is a robust local search-based technique, as the number of parameters increases, convergence can take longer and can get stuck in local minima. Therefore, we are studying the application of a combined DRL-ES controller that can learn and quickly approximately optimize such systems with large numbers of parameters, and then robust ES can keep things stable even as they drift with time.
@@ III-A Simulator
where $\kappa(z,t)\coloneq G(z;u(t))/(\beta\rho)$ and $X^{\prime\prime},Y^{\prime\prime}$ represent $\tfrac{d^{2}X}{dz^{2}},\tfrac{d^{2}Y}{dz^{2}}$. The $\kappa$ terms encode linear focusing/defocusing and the $K$ term models space–charge defocusing. An example solution of the KV equations for the LANSCE LEBT at time $t=t_{0}$ is shown in Fig.2.
The LANSCE LEBT contains $N=22$ independently driven quadrupole magnets. We map the controller’s action vector $u\in\mathbb{R}^{22}$ (one element per magnet) to a piecewise-constant profile as
@@
Figure 3: Architecture of the ES–DRL controller for accelerator tuning. A supervisor selects $\beta$ based on safety constraints and combines $u=\beta(o(t))u_{\mathrm{RL}}+(1-\beta(o(t)))u_{\mathrm{ES}}$. ES may be warm-started from DRL (dotted). The critic is not used during evaluation.
@@ III-B DRL Training Procedure
Algorithm DDPG (off-policy, deterministic)
State $(s)$ dimension $4\times 4000$
Action $(a)$ dimension $22$
Discount $\gamma$ $0.99$
Actor learning rate $1\times 10^{-5}$ (Adam)
Critic learning rate $1\times 10^{-4}$ (Adam)
Replay buffer $|\mathcal{D}|$ $10^{6}$ transitions (uniform)
Batch size $B$ $128$
Target update $\tau$ $0.005$ (soft, every step)
Exploration noise $\mathcal{N}(0,0.1)$ per action
@@
If the IVP solver fails to return valid envelopes in an episode, we assign a large penalty and terminate the episode. The environment is then reset to the last feasible setting. Directly training a controller over all $22$ quadrupole inputs caused frequent failures of the initial–value problem (IVP) solver for certain magnet combinations, which prevented the simulator from returning a solution to the KV equations. To stabilize learning while preserving the intended control objective, we adopt a curricular procedure that progressively increases the control dimension.
@@
Runtime policy: We deploy the policy trained by DDPG. At run time we use only the frozen actor $\mu_{\theta}$ and define the RL control command as
@@ IV Combined ES-DRL Controller
The ES controller is instead uses a finite difference approximation of Eq. ((6)):
@@
$Q_{i}(t+1)=\underbrace{Q_{i}(t)+\Delta t\,\sqrt{\alpha\omega_{i}}\,\cos\!\bigl(\omega_{i}t\Delta t-k\,V(o(t))\bigr)}_{u_{ES,i}},$ (33)
@@
Safety supervisor: The supervisor generates the binary switch $\beta(o(t))$ from envelope measurements. Beam loss increases as the beam envelope approaches the beam–pipe aperture (maximum radius $r_{\max}$), we enforce a safety margin by constraining the beam to remain within $70\%$ of $r_{\max}$. Given the RL action $u_{\mathrm{RL}}$, we integrate the KV equations (17)–(18) to obtain $X(z,t)$ and $Y(z,t)$ along the line (cf. Fig. 2).
@@
$\beta\coloneq\begin{cases}1,&\text{if }\overline{X}(t)\ \text{or}\ \overline{Y}(t)<0.7\,r_{\max},\\[5.69054pt] 0,&\text{otherwise,}\end{cases}$ (34)
so that RL is used when the average envelopes are well within the aperture and control reverts to the robust ES component when either envelope approaches the aperture, which indicates higher beam loss. Other beam-loss models may also be used; see [40]. Applying this rule in (35) switches to ES whenever the RL policy violates the constraint or fails to stabilize the plant. For large $\omega_{i}$, when $\beta=0$ and the ES controller is activated, the ES weak-limit averaging results hold and the overall ES-DRL magnet field dynamics evolve according to
@@
a local model-independent gradient ascent of the time-varying $V(o(t))$, relative to the controlled subset of quadrupole magnets $Q$.
V Simulation Results
@@
where $L_{9-10}^{\star}=176\,$\mathrm{m}\mathrm{m}$\$ is the distance between $Q_{9}$ and $Q_{10}$. The inter-magnet spacing $\Delta L_{9\text{--}10}(t)$ is ramped from 0 to $150\,$\mathrm{m}\mathrm{m}$\$ over steps t=100:200t=100{:}200, held at $0.15\,$\mathrm{m}$\$ for t=200:300t=200{:}300, and ramped back to 0 over t=300:400t=300{:}400, as shown in the top part of Fig. 4. Aggregate performance under these perturbations is summarized in the bottom part of Fig. 4.
@@
VI Conclusions
@@
We presented a hybrid ES–DRL control framework that leverages the complementary strengths of deep reinforcement learning and bounded extremum seeking for stabilizing and optimizing nonlinear time-varying systems. Numerical studies and accelerator-inspired simulations demonstrated that while DRL policies can achieve rapid convergence in-distribution regimes, their performance degrades under unmodeled dynamics and distribution shifts. In contrast, bounded ES guarantees robustness to unknown and drifting control directions, though at the cost of slower convergence. By combining these approaches through a safety-aware supervisor and warm-starting ES from DRL actions, the proposed controller maintained high rewards under severe perturbations and outperformed either method alone. These results suggest a principled path toward deploying learning-based controllers in high-dimensional, safety-critical applications such as particle accelerators, where both adaptability and robustness are essential. Future work will extend this framework to hardware-in-the-loop studies at LANSCE and will include a theoretical study of the convergence properties of this approach for a wide range of nonlinear time-varying dynamic systems.
Acknowledgments
This research has been supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project “Charged Particle Beam Control and Diagnostics using Adaptive Machine Learning,” project number 20220074D. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).

### v2

@@ Abstract
In this paper, we study the use of robust model-independent bounded extremum seeking (ES) feedback control to improve the robustness of deep reinforcement learning (DRL) controllers for a class of nonlinear time-varying systems. DRL has the potential to learn from large datasets to quickly control or optimize the outputs of many-parameter systems, but its performance degrades catastrophically when the system model changes rapidly over time. Bounded ES can handle time-varying systems with unknown control directions, but its convergence speed slows down as the number of tuned parameters increases and, like all local adaptive methods, it can get stuck in local minima. We demonstrate that together, DRL and bounded ES result in a hybrid controller whose performance exceeds the sum of its parts with DRL taking advantage of historical data to learn how to quickly control a many-parameter system to a desired setpoint while bounded ES ensures its robustness to time variations. We demonstrate the generality of our combined ES-DRL controller approach with numerical studies of three very different dynamic systems: 1) a general time-varying system, 2) automatic tuning of the Low Energy Beam Transport section at the Los Alamos Neutron Science Center linear particle accelerator, and 3) an intermittent-contact robotic block pushing task with a time-varying goal.
@@ I Introduction
Deep reinforcement learning (DRL) combines aspects of optimal control theory with deep learning techniques. In DRL, deep neural networks parameterize the policy (controller), the value function, or both. DRL is based on dynamic programming, introduced by Bellman to solve sequential decision problems through the Bellman optimality principle [4]. Classical DP assumes known analytic models of the system dynamics and the reward function, which enables the synthesis of controllers $u$ that maximize cumulative return. Reinforcement learning (RL) relaxes this modeling requirement by learning from data collected as state, action, and reward transitions when the dynamics are unknown [35]. DRL extends RL to high-dimensional state and action spaces by using deep neural networks to approximate value functions and/or policies. This enables the application of RL to complex systems such as image-based perception and continuous control, where tabular methods become intractable [2]. A landmark example is deep Q-learning, which approximates the optimal action value function that maximizes cumulative reward [22]. DRL is being actively investigated for a broad set of applications, including robotic control [36], particle accelerator tuning [15], and the fine-tuning of deep neural networks [24].
@@
In contrast to DRL, the field of control theory has a long history of developing model-independent feedback control algorithms for unknown time-varying systems [37, 17]. For systems of the form
@@
Bounded ES is a new form of ES which has a major advantage of guaranteed bounds on control efforts and parameter update rates despite acting on noisy and analytically unknown and time-varying systems [32]. It has been studied for a wide range of dynamic systems [30], and a general weak-limit averaging analysis has extended its use with non-differentiable dithering functions as well as non-periodic time-varying systems [33]. The ability to stabilize time-varying systems with guaranteed bounds on control effort has made bounded ES useful for safe hardware implementation on high energy systems, where abrupt parameter changes can easily result in damage, such as high energy charged particle accelerators [27, 28]. The bounded ES method has also been applied for GPS-denied source localization [11], optimized path tracking in robotics [3], for biology-inspired 3D source seeking [1], and for tokamak stabilization [7].
@@
While DRL is a powerful method for data-based learning of controllers for dynamic systems, one of the fundamental challenges is handling unknown and rapidly time-varying systems. On the other hand, while adaptive methods such as bounded ES are robust for time-varying systems, they are inherently local feedback-based schemes that do not exploit trajectory histories, are subject to suboptimal solutions, and can converge slowly in high-dimensional parameter spaces. This paper proposes a hybrid framework that combines DRL with bounded ES to leverage the strengths of both. We train a DRL policy using large datasets to find solutions in very few steps when the test dynamics are close to the training distribution. Crucially, the ES layer is warm-started from the DRL policy seeded with DRL-recommended control parameters, which reduces transients and accelerates adaptation when conditions drift. If the system starts to quickly change with time, and the learned DRL policy is no longer valid, robust bounded ES takes over and prevents the system’s performance from severe degradation.
We demonstrate our results with numerical studies of general time-varying dynamic systems, a detailed simulation study of a particle accelerator application with a time-varying magnetic lattice, mimicking natural accelerator behavior, and a detailed simulation study of an intermittent-contact robotic block-pushing task that pushes an object to a time-varying goal position.
@@ II-A Bounded Extremum Seeking for Time-Varying Systems
We briefly recall the results that summarize bounded ES as developed in [32, 30, 33]. For $x\in\mathbb{R}^{n}$, consider system
@@ II-B Deep Reinforcement Learning for Feedback Control
We model DRL as a discounted Markov decision process $M=(S,A,p,r,\gamma)$ with $\gamma\in[0,1)$. At time $t$, the agent observes $s_{t}\!\in\!S$, applies a continuous control $a_{t}\!\in\!A$, receives $r_{t}=r(s_{t},a_{t})$, and transitions to $s_{t+1}\!\sim\!p(\cdot\,|\,s_{t},a_{t})$, with objective
@@ II-C Motivating Example of a Time Varying System
The averaged dynamics control direction is no longer unknown because $b^{2}\geq 0$ and therefore sufficiently large $k\alpha>0$ ensures a gradient ascent of $V(\overline{x})$. As detailed in the bounded ES references, $\omega$ must be chosen sufficiently large relative to $f$. While trivial for bounded ES, this problem is extremely difficult for DRL as shown in Fig. 1.
@@
III ES-DRL for Particle Accelerators
Large particle accelerators are intrinsically time-varying; they are composed of thousands of coupled electromagnetic components whose characteristics drift with temperature, usage, and maintenance; diagnostics are limited and noisy, and they require continual manual retuning. The amplitude and phase of the resonant radio frequency (RF) accelerating cavities drift with time due to temperature, amplifier/distribution effects, and magnet power-supplies ripple and hysteresis break magnetic–field repeatability. Such time-varying behavior is documented for the kilometer-long Los Alamos Neutron Science Center (LANSCE) linear accelerator, for which adaptive ES methods have been developed for beam loss minimization [31]. Because ES is a robust local search-based technique, as the number of parameters increases, convergence can take longer and can get stuck in local minima. Therefore, we are studying the application of a combined ES-DRL controller that can learn and quickly approximately optimize such systems with large numbers of parameters, and then robust ES can keep things stable even as they drift with time.
@@ III-A Simulator
where $\kappa(z,t)\coloneq G(z;u(t))/(\beta\rho)$ and $X^{\prime\prime},Y^{\prime\prime}$ represent $\tfrac{d^{2}X}{dz^{2}},\tfrac{d^{2}Y}{dz^{2}}$. The $\kappa$ terms encode linear focusing/defocusing and the $K$ term models space–charge defocusing. An example solution of the KV equations for the LANSCE LEBT at time $t=t_{0}$ is shown in Fig. 2. The LANSCE LEBT contains $N=22$ independently driven quadrupole magnets. We map the controller’s action vector $u\in\mathbb{R}^{22}$ (one element per magnet) to a piecewise-constant profile as
@@
Figure 3: Architecture of the ES–DRL controller for accelerator tuning. A supervisor selects binary $\beta\in\{0,1\}$ based on safety constraints and combines $u=\beta(o(t))u_{\mathrm{RL}}+(1-\beta(o(t)))u_{\mathrm{ES}}$. ES may be warm-started from DRL (dotted).
@@ III-B DRL Training Procedure
Hyperparameter Particle Accelerator Robot
State $(s)$ dimension $4\times 4000$ 28
Action $(a)$ dimension $22$ 4
Discount $\gamma$ $0.99$ 0.99
Actor learning rate $\times 10^{-5} 1×1$$\times 10^{-4}$
Critic learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
Replay buffer $|\mathcal{D}|$ $10^{6}$ transitions (uniform) $10^{6}$ transitions
Batch size $B$ $128$ 256
Target update $\tau$ $0.005$ (soft, every step) 0.005
Exploration noise $\mathcal{N}(0,0.1)$ per action $\mathcal{N}(0,0.1)$
@@
If the IVP solver fails to return valid envelopes we assign a large penalty and terminate the episode. The environment is then reset to the last feasible setting to continue exploration from a known-good configuration near the constraint boundary rather than repeatedly returning to the initial operating point. Directly training a controller over all $22$ quadrupole inputs caused frequent failures of the initial–value problem (IVP) solver for certain magnet combinations, which prevented the simulator from returning a solution to the KV equations. To stabilize learning while preserving the intended control objective, we adopt a curricular procedure that progressively increases the control dimension.
@@
Runtime policy: We deploy the policy trained by DDPG. At run time we use only the frozen actor $\mu_{\theta}$ to prevent the policy drift and define the RL control command as
@@ III-C Combined ES-DRL Controller
The ES controller instead uses a finite difference approximation of Eq. (6):
@@
$Q_{i}(t+1)=\underbrace{Q_{i}(t)+\Delta t\,\sqrt{\alpha\omega_{i}}\,\cos\!\bigl(\omega_{i}t\Delta t-k\,V(o(t))\bigr)}_{u_{ES,i}}.$ (33)
@@
Safety supervisor: The supervisor generates the binary switch $\beta(o(t))$ from envelope measurements. It is evaluated at every discrete control step, so the handoff occurs immediately once the envelope threshold is violated. Beam loss increases as the beam envelope approaches the beam–pipe aperture; we enforce a safety margin by constraining the beam to remain within $70\%$ of $r_{\max}$. Given the RL action $u_{\mathrm{RL}}$, we integrate the KV equations (17)–(18) to obtain $X(z,t)$ and $Y(z,t)$ along the line (cf. Fig. 2).
@@
$\beta\coloneq\begin{cases}1,&\text{if }\overline{X}(t)\ \text{or}\ \overline{Y}(t)<0.7\,r_{\max}\text{(RL mode)},\\[5.69054pt] 0,&\text{otherwise (ES mode),}\end{cases}$ (34)
so that RL is used when the average envelopes are well within the aperture and control reverts to the robust ES component when either envelope approaches the aperture, which indicates higher beam loss. Other beam-loss models may also be used [38]. Applying this rule in Eq.(30) switches to ES whenever the RL policy violates the constraint or fails to stabilize the plant. The ES weak-limit averaging results hold and the overall ES-DRL magnet field dynamics evolve according to
@@
a local model-independent gradient ascent of the time-varying $V(o(t))$, relative to the controlled subset of $Q$.
III-D Simulation Results
@@
where $L_{9-10}^{\star}=176\,$\mathrm{m}\mathrm{m}$\$ is the distance between $Q_{9}$ and $Q_{10}$. The inter-magnet spacing $\Delta L_{9\text{--}10}(t)$ is ramped from 0 to $150\,$\mathrm{m}\mathrm{m}$\$ over steps t=100:200t=100{:}200, held at $0.15\,$\mathrm{m}$\$ for t=200:300t=200{:}300, and ramped back to 0 over t=300:400t=300{:}400, as shown in the top part of Fig. 4. Aggregate performance under these perturbations is summarized in Fig. 4.
@@
IV ES-DRL controller for Intermittent Contact robot task with Time-Varying Goal
Next, we apply the same ES-DRL controller in (30) to the control of a robotic arm, as shown in Fig. 5. The task consists of a 7-DoF Fetch mobile manipulator arm that must push a movable block on a tabletop to a desired goal position in the FetchPush benchmark environment [25]. Robotic block pushing is a widely studied manipulation task, with recent work demonstrating robust pushing using force feedback alone [14]. The desired goal follows a smooth time-varying trajectory in the tabletop plane: at each step $t$ it traces a circular path of radius $0.10$ m about a nominal center $(g_{x0},g_{y0})$ with a period of $200$ steps, i.e., $g_{x}(t)=g_{x0}+0.10\sin(2\pi t/200)$ and $g_{y}(t)=g_{y0}+0.10\cos(2\pi t/200)$, and $g_{z}(t)$ is fixed, which produces a distribution shift relative to the stationary-goal settings used for RL training. The observation is a 25-dimensional vector containing the end-effector Cartesian position and velocity, the block position, the relative block-to-gripper position, the left and right gripper finger joint displacements and velocities, the block XYZ Euler orientation, and the block linear and angular velocities. In our implementation we concatenate observation and desired_ goal to form a state vector $s_{t}\in\mathbb{R}^{28}$. Actions are continuous Cartesian displacement commands for the end-effector together with a gripper command $g(t)$,
$a_{t}\in\mathbb{R}^{4},\qquad a_{t}=[\Delta x(t),\,\Delta y(t),\,\Delta z(t),\,g(t)]^{\top},$ (36)
which are applied through the MuJoCo simulator’s internal mocap operational-space controller to generate the low-level joint actuation. The DDPG training hyperparameters are summarized in Table I. During evaluation, exploration noise is disabled and only the trained actor is used to compute the RL actions.
Reward design: We use a simple dense shaping reward based on two Euclidean distances: $d_{1}$, between the end-effector and the block, and $d_{2}$, between the block and the goal. At each step we set $r_{t}=-(d_{1}+d_{2})$, and add a success bonus of +2+2 when the block reaches the goal.
Safety supervisor: For ES-DRL controller, we use RL for fast approach to the goal and switch to bounded ES once physical interaction begins. Let $c_{t}\in\{0,1\}$ be a contact flag that indicates whether the end-effector is in contact with the block at time $t$. We define a supervisor
$\beta_{t}\in\{0,1\},\qquad\beta_{t}=\begin{cases}1,&c_{t}=0\quad(\text{RL mode}),\\ 0,&c_{t}=1\quad(\text{ES mode}).\end{cases}$ (37)
Results: Fig. 5 highlights the effect of goal drift as an out-of-distribution disturbance for the learned policy. Although the RL controller initially approaches and begins to push the block, the time-varying target drives the policy outside the training distribution; the gripper subsequently loses effective contact and the block stagnates away from the moving goal. ES is more robust to this nonstationarity and eventually discovers a pushing direction, but it requires a longer, more exploratory approach to first acquire contact and then align the push, resulting in a larger path length. In contrast, ES-DRL leverages RL for rapid and directed approach to establish contact, then switches to ES during interaction to adapt the push online, preserving contact and reaching the time-varying goal more quickly and with a more direct trajectory.
Figure 5: Time-varying goal trajectory and block position.
V Conclusions
@@
We presented a hybrid ES–DRL control framework that leverages the complementary strengths of deep reinforcement learning and bounded extremum seeking for stabilizing and optimizing nonlinear time-varying systems. Numerical studies of general time-varying systems, particle accelerator beams, and robot arms demonstrated that while DRL policies can achieve rapid convergence in-distribution regimes, their performance degrades under unmodeled dynamics and distribution shifts. In contrast, bounded ES guarantees robustness to unknown and drifting control directions, though at the cost of slower convergence. By combining these approaches through a safety-aware supervisor and warm-starting ES from DRL actions, the proposed controller maintained high rewards under severe perturbations and outperformed either method alone. These results suggest a principled path toward deploying learning-based controllers in high-dimensional, safety-critical applications such as particle accelerators and robotic arms, where adaptability and robustness are essential.